# APISR: Anime Production Inspired Real-World Anime Super-Resolution

### v1

@@ Abstract
While real-world anime super-resolution (SR) has gained increasing attention in the SR community, existing methods still adopt techniques from the photorealistic domain. In this paper, we analyze the anime production workflow and rethink how to use characteristics of it for the sake of the real-world anime SR. First, we argue that video networks and datasets are not necessary for anime SR due to the repetition use of hand-drawing frames. Instead, we propose an anime image collection pipeline by choosing the least compressed and the most informative frames from the video sources. Based on this pipeline, we introduce the Anime Production-oriented Image (API) dataset. In addition, we identify two anime-specific challenges of distorted and faint hand-drawn lines and unwanted color artifacts. We address the first issue by introducing a prediction-oriented compression module in the image degradation model and a pseudo-ground truth preparation with enhanced hand-drawn lines. In addition, we introduce the balanced twin perceptual loss combining both anime and photorealistic high-level features to mitigate unwanted color artifacts and increase visual clarity. We evaluate our method through extensive experiments on the public benchmark, showing our method outperforms state-of-the-art approaches by a large margin. The code is available at https://github.com/Kiteretsu77/APISR.
footnotetext: † Corresponding author.
@@ 2 Realated Work
Anime represents a distinctive form of visual art, often characterized by exaggerated visual representation. Creators of anime typically start by sketching line art, followed by 2D and 3D animation techniques, which include elements like colorization, CGI effects, and frame interpolation. Notably, recent research in the realm of anime has garnered substantial attention, e.g., AI painting with anime content [59, 15], vectorization of anime images [60, 55], anime interpolation and inbetweening [34, 9, 36], anime sketch colorization [5, 4, 43, 58], 3D representation [35, 10], and anime domain adaptation [21].
@@ 4.2 Comparisons with State-of-the-art Methods
We compare our APISR quantitatively and qualitatively with other SOTA real-world image and video SR methods, which includes Real-ESRGAN [46], BSRGAN [57], Real-BasicVSR [7], AnimeSR [48], and VQD-SR [39].
Table 1: Quantitative comparisons on AVC-RealLQ[48]. Bold text indicates the best performance. (‘***’ denotes fine-tune on animation videos from  [48])
@@ 4.3 Ablation Study
As shown in Tab. 2, we substitute our API training dataset with several alternatives for comparative analysis: AVC-Train [48], frames randomly selected from the same video source as our API, a collection of I-Frames with IQA selection, and a collection of I-Frames with ICA selection. For a fair comparison, we keep the same training dataset size. If we take the AVC-Train video training dataset as an image dataset to train, we include temporal distorted images and less informative frames, which makes the performance hard to compete with the model trained with API in all metrics. Randomly selected image datasets perform poorly because they lack attention to high-quality frames in videos. With our I-Frame collection, we take off temporally distorted frames and choose the least compressed frames, but IQA-based selection limits the performance. With the same training iterations and conditions, the dataset selected by ICA-based criteria leads to an improvement over the dataset by IQA-based selection. With the 720P rescaling method, anime images have more compact hand-drawn lines and CGI information than falsely upscaled versions, and this back-to-original thinking boosts the performance in all metrics.

### v2

@@ Abstract
While real-world anime super-resolution (SR) has gained increasing attention in the SR community, existing methods still adopt techniques from the photorealistic domain. In this paper, we analyze the anime production workflow and rethink how to use characteristics of it for the sake of the real-world anime SR. First, we argue that video networks and datasets are not necessary for anime SR due to the repetition use of hand-drawing frames. Instead, we propose an anime image collection pipeline by choosing the least compressed and the most informative frames from the video sources. Based on this pipeline, we introduce the Anime Production-oriented Image (API) dataset. In addition, we identify two anime-specific challenges of distorted and faint hand-drawn lines and unwanted color artifacts. We address the first issue by introducing a prediction-oriented compression module in the image degradation model and a pseudo-ground truth preparation with enhanced hand-drawn lines. In addition, we introduce the balanced twin perceptual loss combining both anime and photorealistic high-level features to mitigate unwanted color artifacts and increase visual clarity. We evaluate our method through extensive experiments on the public benchmark, showing our method outperforms state-of-the-art anime dataset-trained approaches. The code is available at https://github.com/Kiteretsu77/APISR.
footnotetext: † Corresponding authorfootnotetext: * works done at University of Michigan
@@ 2 Realated Work
Anime represents a distinctive form of visual art, often characterized by exaggerated visual representation. Creators of anime typically start by sketching line art, followed by 2D and 3D animation techniques, which include elements like colorization, CGI effects, and frame interpolation. Notably, recent research in the realm of anime has garnered substantial attention, e.g., AI painting with anime content [67, 18, 22], vectorization of anime images [68, 63], anime interpolation and inbetweening [40, 10, 42], anime sketch colorization [6, 5, 50, 66, 13], 3D representation [41, 11], and anime domain adaptation [26].
@@ 4.2 Comparisons with State-of-the-art Methods
We compare our APISR quantitatively and qualitatively with other SOTA real-world image and video SR methods, which include Real-ESRGAN [54], BSRGAN [65], Real-BasicVSR [8], AnimeSR [56], and VQD-SR [46].
Table 1: Quantitative comparisons on AVC-RealLQ [56]. Bold text indicates the best performance. (‘*’ denotes fine-tune on animation videos from  [56])
@@ 4.3 Ablation Study
As shown in Tab. 2, we substitute our API training dataset with several alternatives for comparative analysis: AVC-Train [56], frames randomly selected from the same video source as our API, a collection of I-Frames with IQA selection, and a collection of I-Frames with ICA selection. For a fair comparison, we keep a similar intensity of the training dataset size. If we take the AVC-Train video training dataset as an image dataset to train, we include temporal distorted images and less informative frames, which makes the performance hard to compete with the model trained with API in all metrics. Randomly selected image datasets perform poorly because they lack attention to high-quality frames in videos. With our I-Frame collection, we take off temporally distorted frames and choose the least compressed frames, but IQA-based selection limits the performance. With the same training iterations and conditions, the dataset selected by ICA-based criteria leads to an improvement over the dataset by IQA-based selection. With the 720P rescaling method, anime images have more compact hand-drawn lines and CGI information than falsely upscaled versions, and this back-to-original thinking boosts the performance in all metrics.
@@ 5 Conclusion
Supplementary Material
In this supplementary material, Sec. A first presents more statistics and details of our proposed anime image SR training dataset. Then, Sec. B shows details about our implementations in super-resolution (SR) network training. Specifically, Sec. B.1 presents the image SR network we used in our training. Sec. B.2 presents details of post-processing techniques we use on the pseudo-GT preparation for hand-drawn line enhancement. Sec. B.3 presents figures and details of the ResNet50 [21] perceptual loss for our proposed balanced twin perceptual loss. Sec. B.4 provides the hyperparameter setting for our proposed prediction-oriented compression and shuffled resize module in the degradation model. Finally, Sec. C provides more visual results of comparisons among SOTA methods and ablation studies.
Appendix A API Dataset Details
Our Anime Production-oriented Image (API) SR dataset contains 3,740 high-quality and informative images. This quantity is roughly the same quantity as the previous photorealistic SR training dataset size [54, 65], which includes DIV2K [1], Flickr2K [45], and OutdoorSceneTraining [52]. The aspect ratio and resolution information before scaling are shown in Fig. 11.
Appendix B Implementation Details
B.1 Training Network Details
The generator network we deploy is GRL [30], a SOTA image SR network (CVPR 2023). GRL leverages interconnected relationships within various layers of image structures through a Transformer-based framework, attaining improvement in multiple tasks of SR and image restoration. The model we chose is its tiny version, which has 0.91M parameters. To better adapt the real-world SR task, we changed its upsampler module from the default pixel shuffle strategy to the nearest neighbor interpolation with the convolution layer approach, which is used for the base model version but not for the tiny version in their proposed methods. We change the upsampler because the nearest neighbor interpolation with the convolution layer is claimed to show fewer artifacts in the upsampling process than the pixel shuffle strategy. The final network parameter is 1.03M, which is the smallest network among all image and video-based SOTA methods that we compare.
B.2 Hand-drawn line enhancement Details
In the hand-drawn line enhancement, we have proposed outlier filter and passive dilate techniques to obtain a clean XDoG-extracted [55] hand-drawn line edge map. XDoG is widely used in paired dataset preparation in anime colorization [6, 5, 50, 22]. The extracted edge map by XDoG is a binary output, where the white pixel stands for the active edge map region and the black pixel stands for the unrelated region.
For the outlier filter, we use breadth-first search in eight directions to recursively detect the surrounding pixels of all white pixels and turn white pixel regions into black pixels if the total quantity of connected white pixels is less than the threshold. We empirically set the threshold as 32.
For the dilation, we passively replace the black pixel with the white pixel if it has more than 3 white pixel neighbors, which is different from independent kernel-based active dilation methods in [29, 12, 20] that directly spread the surrounding neighbors to be white pixels if the central pixel is white. Compared to active dilation methods, our proposed passive dilation is more concentrated on the hand-drawn lines region instead of covering unrelated pixel information (see Fig. 13). Thus, we name our methods as passive dilatation.
In the implementation, we will do an unsharp mask for the whole image first to increase overall visualization sharpness and then apply two extra turns of sharpening to the hand-drawn lines specifically based on the pipeline design mentioned above. More implementation details can be found in our released code.
B.3 Balanced Twin Perceptual Loss Details
As shown in Fig. 12, our proposed middle-layer output comparisons for ResNet50 [21] follow the idea proposed by ESRGAN [53] which compares feature map outputs before the activation layer. Following VGG-based perceptual loss [27], we compare the last convolution layer of each stage. There are five middle-layer output comparisons, which are the same quantity as VGG-based perceptual loss [27]. Thus, our proposed twin perpetual loss reaches a mutual balance in training.
Figure 11: API dataset extra statistics.
Figure 12: The overview of our proposed middle-layer outputs of ResNet50 [21] perceptual loss trained by Danbooru dataset [3]. Overall, ResNet50 can be summarized into five stages which is similar to VGG [39]. $\phi_{j}$ represents the perceptual function that returns $j$th layer output of ResNet50.
Figure 13: Comparisons between active and passive dilation. Our proposed passive dilation is more concentrated on the hand-drawn line region without producing over-sharpened pseudo-GT images as in active dilation methods.
B.4 Degradation Details
For the prediction-oriented compression module of the degradation model, we deploy both the image compression with prediction mechanism (i.e., WebP [38] and AVIF [19]) and single-frame video compression. Meanwhile, for the robustness of the degradation model, we keep the JPEG [47]. The quality factor range of JPEG, WebP, and AVIF is $[20,95]$ with encoding speed in the range of $[0,6]$ for WebP and AVIF. The probability of fetching the value in the range is equal.
For the stability of video compression processing, we choose the widely-used video processing tools, ffmpeg, to perform the proposed single-frame compression of MPEG2 [32], MPEG4 [2], H.264 [36], and H.265 [44]. In ffmpeg, CRF is an engineering system to control the quantization level, and preset is a speed control mechanism whose setting is directly related to compression distortions. For MPEG2 and MPEG4, we empirically find that the quality factor control (-qscale:v) is a better way to control single-frame compression, but for H.264 and H.265, CRF is a better way to control. For MPEG2 and MPEG4, we set the quality factor in the range $[8,31]$. For H.264 and H.265, we set the CRF in the range $[23,38]$ and $[28,42]$ respectively. The preset for all of them is $\{slow,medium,fast,faster,superfast\}$ with probability $\{0.05,0.35,0.3,0.2,0.1\}$.
The first prediction-oriented compression includes JPEG [47] and WebP [38] with a probability of $\{0.4,0.6\}$ respectively. The second prediction-oriented compression includes JPEG [47], WebP [38], AVIF [19], and single-frame compression of MPEG2 [32], MPEG4 [2], H.264 [36], and H.265 [44] with probability of $\{0.06,0.1,0.1,0.12,0.12,0.3,0.2\}$ respectively. For the first resize module, we set the scaling in the range of $[0.1,1.2]$ with probability $\{0.2,0.7,0.1\}$ to scale up, scale down, or remain current resolution. For the second resize module, we choose the range of $[0.15,1.2]$ with probability $\{0.2,0.7,0.1\}$. More implementation details can be found in our released code.
Appendix C More Qualitative Comparisons
In this section, we present more qualitative results to verify the effectiveness of our APISR among SOTA methods. Moreover, we provide visual comparisons for the ablation studies.
Extra Qualitative Comparisons with SOTA methods.
Fig. 14 and Fig. 15 show extra qualitative comparisons on AVC-RealLQ [56] datasets for 4×4$\times$4 × scaling. This includes image-based Real-ESRGAN [54] and BSRGAN [65], and video-based RealBasicVSR [8], AnimeSR [56], and VQD-SR [46]. Our APISR presents clearer and sharper hand-drawn lines (first example of Fig. 14, first and second examples of Fig. 15, and third example of Fig. 16), better restoration with more natural details (second and third examples of Fig. 14, and third example of Fig. 15), and does not present unwanted color artifacts (first and second examples of Fig. 16).
Qualitative Comparisons of Ablation Studies.
Fig. 17, Fig. 18, and Fig. 19 shows the qualitative comparisons of ablations studies.
As shown in Fig. 17, the network trained with AVC-Train [56] over-sharpens the grid texture and produces annoying artifacts as denoted by the arrows in the figure. Similarly, the network trained with the random sampled or IQA-based sampled dataset can alleviate this artifact but is still hard to completely remove it. However, when we introduce the ICA-based selection method with I-Frame dataset collection, this artifact is greatly removed and the generated image shows more natural details. This is thanks to versatile complex scenes included in the dataset due to ICA-based selection. With 720P rescaling, fewer ringing artifacts appear.
As shown in Fig. 18, the network trained with high-order [54] and random order [65] degradation model presents ringing artifacts, rainbow effects, and color distortions as denoted by the arrows in the figure. Nevertheless, introducing our proposed prediction-oriented compression module in the degradation model promotes the network to greatly restore these problems and generate more natural details with less distorted hand-drawn lines. Moreover, with the shuffled resize module in the degradation model, more distortions are restored and present natural shadow details.
As shown in Fig. 19, the network trained with the plain version presents unwanted color pixel artifacts and sparse hand-drawn line information as denoted by the arrows in the figure. With the hand-drawn line enhancement, the hand-drawn line around the eyes of the character is greatly intensified and more details are generated. However, the unwanted color pixels still exist and they are presented as an annoying artifact. With the twin perceptual loss, the unwanted color pixels are greatly alleviated. Further, with the scaling to early layers in ResNet perceptual loss, more shadow artifacts and distortions are restored.
Figure 14: Qualitative comparisons on AVC-RealLQ [56] for 4×4$\times$4 × scaling. Our APISR presents clearer and sharper hand-drawn lines, better restoration with more natural details, and does not present unwanted color artifacts. Zoom in for the best view.
Figure 15: Qualitative comparisons on AVC-RealLQ [56] for 4×4$\times$4 × scaling. Our APISR presents clearer and sharper hand-drawn lines, better restoration with more natural details, and does not present unwanted color artifacts. Zoom in for the best view.
Figure 16: Qualitative comparisons on AVC-RealLQ [56] for 4×4$\times$4 × scaling. Our APISR presents clearer and sharper hand-drawn lines, better restoration with more natural details, and does not present unwanted color artifacts. Zoom in for the best view.
Figure 17: Qualitative comparisons of the first ablation study. IQA stands for image quality assessment. ICA stands for image complexity assessment. 720P stands for our proposed 720P rescaling. Zoom in for the best view.
Figure 18: Qualitative comparisons of the second ablation study. Zoom in for the best view.
Figure 19: Qualitative comparisons of the third ablation study. Hand-drawn lines enhancement is denoted as Sharpen and twin perceptual loss is denoted as APL. Balanced Scale presents the early layer scaling to ResNet perceptual loss. Zoom in for the best view.